# Stable generative modeling using diffusion maps

### v1

@@ Abstract.
We consider the problem of sampling from an unknown distribution for which only a sufficiently large number of training samples are available. Such settings have recently drawn considerable interest in the context of generative modelling. In this paper, we propose a generative model combining diffusion maps and Langevin dynamics. Diffusion maps are used to approximate the drift term from the available training samples, which is then implemented in a discrete-time Langevin sampler to generate new samples. By setting the kernel bandwidth to match the time step size used in the unadjusted Langevin algorithm, our method effectively circumvents any stability issues typically associated with time-stepping stiff stochastic differential equations. More precisely, we introduce a novel split-step scheme, ensuring that the generated samples remain within the convex hull of the training samples. Our framework can be naturally extended to generate conditional samples. We demonstrate the performance of our proposed scheme through experiments on synthetic datasets with increasing dimensions and on a stochastic subgrid-scale parametrization conditional sampling problem.
@@ 1. Introduction
Generative modeling is the process of learning a mechanism for synthesizing new samples that resemble those of the original data-generating distribution, given only a finite set of samples. It has seen wide adoption and enormous success across diverse application domains, from image [5, 21, 43, 49] and text generation [63, 29, 30], to drug discovery [3, 2] and anomaly detection [10, 47], to name but a few.
In this paper, we introduce a new nonparametric approach to generative modeling that combines ideas from optimal transportation, diffusion maps, and Langevin dynamics.
Suppose that we are given $M$ training samples $x^{(i)}\sim\pi$, $i=1,\ldots,M$, from an unknown distribution $\pi$ on $\mathbb{R}^{d}$. Perhaps the simplest nonparametric approach to generative modeling is to build a kernel density estimate (KDE) and then sample from it; the KDE is essentially a mixture model with $M$ components. Alternatively, one could estimate the score function, $\hat{s}_{M}(x)\approx s(x)\coloneqq\nabla\log\pi(x)$, without directly estimating $\pi$, and use this estimate as the drift term of Langevin dynamics,
(1) $\dot{X}_{t}=\hat{s}_{M}(X_{t})+\sqrt{2}\dot{W}_{t},$
where $W_{t}$ denotes standard $d$-dimensional Brownian motion.
There are myriad ways of estimating the score function [22, 54], and given an estimate for it, one needs to discretize (1), for example using Euler–Maruyama, to obtain an implementable scheme. However, the step size needs to be carefully chosen: a small step size leads to slow convergence, while too large a step yields instability of the numerical scheme, especially for data that are supported on a compact manifold, e.g.,
@@
for some unknown function $g(x)$. Any estimated score function $\hat{s}_{M}(x)$ will take large values for $x$ with $\|g(x)\|^{2}\gg 0$, rendering the Langevin dynamics (1) stiff. In other words, an explicit time integration method such as Euler–Maruyama will require extremely small step sizes.
In this paper, we choose an alternative approach. Instead of first estimating the score function $\hat{s}_{M}(x)$ and then discretising (1) in time, we employ diffusion maps [8, 7, 37] to directly approximate the semigroup $\exp(\epsilon\mathcal{L})$, $\epsilon>0$, of a diffusion process with generator $\mathcal{L}$ and invariant distribution $\pi$, from the given samples $\{x^{(i)}\}_{i=1}^{M}$. The second key ingredient of our method is to interpret $\epsilon$ as a step size and to read off a Gaussian transition kernel from the diffusion map in the form of
(3) $X_{n+1}=m_{\epsilon}(X_{n})+\Sigma(X_{n})\Xi_{n},$
@@
with appropriately defined functions $m_{\epsilon}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\Sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$, and $\Xi_{n}\sim{\rm N}(0,I)$. Broadly, $m_{\epsilon}(X_{n})$ controls the drift, while $\Sigma(X_{n})\Xi_{n}$ introduces noise. In this recursion, the step size $\epsilon$ is linked with the error of the diffusion map approximation, making the time discretization implicit. Comparing to directly discretizing (1) using Euler–Maruyama with step-size $\Delta t=\epsilon$, we will demonstrate that the scheme (3) is stable and ergodic for all step-sizes $\epsilon>0$ and, hence, $\epsilon$ can be chosen solely on accuracy considerations. Furthermore, the inclusion of the position-dependent diffusion matrix $\Sigma(x)$ makes it better suited for sampling from a manifold.
@@
While our approach is rooted in diffusion maps [8, 7, 37, 38], our construction utilizes a Sinkhorn scheme [58], which is motivated by considering a special case of the discrete Schrödinger bridge problem of optimally coupling the empirical measure of the training samples with itself. By solving the Schrödinger bridge problem, we construct a transition matrix whose state space encompasses all the training points. We generalize this approach to the continuous state space that extends beyond the current training data points. Armed with this transition kernel, we obtain a Markov chain that samples from the underlying distribution of the training data via (3). From here, we introduce a novel split-step time-stepping scheme, which ensures that the generated samples consistently lie within the convex hull of the training samples. In contrast, using a direct discretization of (1) with Euler–Maruyama results in generating samples on unbounded domains.
@@
In addition, we replace the fixed bandwidth kernel with a variable bandwidth kernel within our diffusion map framework. As we demonstrate in our numerical experiments, the resulting sampling scheme (3) provides a better representation of the underlying target distribution. More precisely, we assess the quality of the generated samples using a variable bandwidth kernel and using a fixed bandwidth kernel on synthetic data sets drawn from non-uniform distributions supported on irregular domains and on low-dimensional manifolds.
@@
We then extend our method to create a conditional generative model. This allows to perform Bayesian inference in the “simulation-based” setting, i.e., without explicit evaluation of a prior density or likelihood. We demonstrate the performance of this approach in a stochastic subgrid-scale parametrization problem.
@@ 1.1. Related work
Langevin dynamics (1) characterizes the motion of particles as they experience a blend of deterministic and stochastic forces. Unlike in this paper, it is typically assumed that the deterministic forcing term $\nabla\log\pi(x)$ is given. Langevin dynamics has been used as a popular tool for sampling data from the target distribution $\pi$. One variation of this is to introduce a symmetric preconditioning operator to the Langevin dynamics. Some popular choices of the preconditioning include the empirical covariance [12, 6, 39] and the Riemannian metric [13, 61, 28], making this method converge faster and more geometry-aware, while leaving the stationary distribution unchanged.
@@
On the other hand, diffusion maps have traditionally served as a tool for nonlinear dimensional reduction [8, 7, 37]. The kernel matrix formed using pairwise sample distances with appropriate normalization approximates the semigroup $\exp(\epsilon\mathcal{L})$. In recent work [58], an alternative normalization technique based on Sinkhorn weights has been studied. The Sinkhorn algorithm solves for the Markov transition kernel associated with a discrete Schrödinger bridge problem, where the coupling is between the empirical measures of the training samples with themselves. This approach results in a symmetric stochastic operator that, notably, also approximates the semigroup $\exp(\epsilon\mathcal{L})$. The drift term of the Langevin dynamics is estimated by acting on the identity function. Separately, the idea of using variable bandwidth kernels can be found very early in the statistics community for kernel density estimation [45, 56], for estimating regression curves [36] and mean regression functions [11]. Recently, [1] replaces the original fixed bandwidth kernel with the variable bandwidth kernel in the construction of diffusion maps, making the approximation of the generator accurate on unbounded domains. Inspired by this concept, we replace the fixed bandwidth kernel with a variable bandwidth kernel. The resulting normalized matrix approximates the semigroup of a different Langevin diffusion process.
@@
In recent years, there has been a surge of research interest in the realm of generative modeling. Despite the able achievements of well-established neural network-based generative models, such as variational auto-encoders (VAE) [25, 42], generative adversarial networks (GAN) [14], and diffusion models or score generative models (SGM) [20, 51, 55, 62], they often require meticulous hyperparameter tuning [44, 53] and exhibit a long training time [14, 57, 50]. The efficacy of these methods significantly hinges on the architectural choices and parameter settings of the underlying deep neural networks [46, 24], which, regrettably, demands a high level of expertise. Furthermore, SGMs solve both a forward and a reverse stochastic differential equation (SDE). The forward SDE introduces noise to the sample, transforming the data into the standard normal distribution, while the reverse SDE takes sample from the standard normal distribution back to the original data distribution, yielding a different sample than the one initially fed into the forward SDE. During the training process, the score function is learned, not for the target distribution, but for the data distribution at each time. Our work, on the other hand, solves only one (forward) SDE, and we learned the score of the target distribution only once.
@@
Several recent studies have combined a range of score function estimation techniques with Langevin dynamics. For example, [52] introduces a noise conditional score network to learn the score function and then uses annealed Langevin dynamics to generate samples, and [4] studies the convergence rate of a Langevin based generative model, where the score is estimated using denoising auto-encoders. Such techniques are also studied within the Bayesian imaging community, commonly referred to as “plug and play” [27]. Nevertheless, these approaches use neural networks for the estimation of the score function, necessitating substantial fine-tuning, and their effectiveness depends on factors such as the complexity of the approximation families and the architectural structures of the neural networks. In addition, [27] uses an explicit projection to ensure that samples stay on the compact manifold given by (2) which is assumed to be explicitly known. In contrast, we do not assume any knowledge of $\mathcal{M}$.
1.2. Outline.
@@
In Section 2 we construct a Markov chain using a Schrödinger bridge diffusion map approximation that samples from the given discrete data distribution. In Section  3, we extend this Markov chain to the continuous setting by constructing a Gaussian transition kernel which extracts its conditional mean and covariance matrix from the underlying diffusion map approximation. We introduce two discrete-time Langevin samplers; one with an arbitrary data-unaware diffusion and one with a data-aware diffusion matrix in Section 3.1. Theoretical properties such as stability and ergodicity are discussed in Section 3.2. We further discuss the application of variable bandwidth kernels when estimating the diffusion map in Section 3.3. While Section 3 focuses on finite step-size and finite sample-size implementations, Section 4 establishes connections to the underlying semi-groups and generators in the infinite sample-size limit. We explore the extension of our proposed scheme to a conditional sampling setting in Section 5, and we demonstrate our proposed methods in Section 6 in a suite of examples, including a conditional sampling exercise with an application to stochastic subgrid-scale parametrization. We conclude in Section 7 with a summary and an outlook.
@@ 2. Discrete Schrödinger bridges
We consider the Schrödinger bridge problem of coupling $\mu_{\rm em}$ with itself along a reversible reference process of the form
(4) $X^{\prime}=X+\sqrt{\epsilon}\left(K(X^{\prime})+K(X)\right)^{1/2}\Xi,$
where $\epsilon>0$ is a tuneable parameter, $\Xi\sim{\rm N}(0,I)$, and $K(x)$ is a symmetric positive definite matrix for all $x\in\mathbb{R}^{d}$. Popular choices include $K=I$, $K=\Sigma_{M}$, where $\Sigma_{M}$ is the empirical covariance matrix of the samples $\{x^{(i)}\}_{i=1}^{M}$, and $K=\rho(x)I$, where $\rho(x)>0$ is a scaling function representing variable bandwidth. The update step (4) corresponds to a Stratonovitch SDE with multiplicative noise and diffusion matrix $K(x)$. This provides a natural way of interpreting $K(X^{\prime})+K(X)$ as a preconditioner for the associated Schrödinger bridge problem, and this fact will be explored in the context of variable bandwidth implementations in Section 3.3.
Instead of working with the empirical measure $\mu_{\rm em}({\rm d}x)$, we introduce the probability vector $p^{\ast}=(1/M,\ldots,1/M)^{\rm T}\in\mathbb{R}^{M}$ over $\{x^{(i)}\}_{i=1}^{M}$. Then the Schrödinger bridge problem is solved by first introducing the symmetric matrix $T_{\epsilon}\in\mathbb{R}^{M\times M}$ with entries
(5) $t_{ij}=\exp\left(-\frac{1}{2\epsilon}(x^{(i)}-x^{(j)})^{\rm T}\left(K(x^{(i)})+K(x^{(j)})\right)^{-1}(x^{(i)}-x^{(j)})\right).$
One then finds the scaling vector $v_{\epsilon}\in\mathbb{R}^{M}$ such that the symmetric matrix
@@
satisfies
$P_{\epsilon}{\rm 1}_{M}=p^{\ast}.$
Here ${\rm 1}_{M}=(1,\ldots,1)^{\rm T}\in\mathbb{R}^{M}$, and $D(v)\in\mathbb{R}^{M\times M}$ denotes the diagonal matrix with diagonal entries provided by $v\in\mathbb{R}^{M}$.
Given $P_{\epsilon}$, one can now construct a Markov chain that samples from $\mu_{\rm em}$. Assume the Markov chain is currently in state $x^{(j)}$, then the transition probabilities to the next state $x\in\{x^{(i)}\}_{i=1}^{M}$ are given by
$p_{j}=MP_{\epsilon}e_{j}\in\mathbb{R}^{M},$
where $e_{j}\in\mathbb{R}^{M}$ denotes the $j$-th unit vector in $\mathbb{R}^{M}$. Since all entries in $P_{\epsilon}$ are bounded from below provided all samples satisfy $x^{(i)}\in\mathcal{M}$, where $\mathcal{M}$ is a compact submanifold in $\mathbb{R}^{d}$, the constructed Markov chain possesses a unique invariant measure given by $p^{\ast}$ and is geometrically ergodic. The rate of convergence can be determined by the diffusion distance
@@
If the diffusion distance is small, then $x^{(i)}$ and $x^{(j)}$ are well connected. Furthermore, if $d(x^{(i)},x^{(j)})$ is small for all points, then the Markov chain will mix quickly. In particular, larger values of $\epsilon$ will lead to faster mixing. This idea can be further extended to the case where the samples are not uniformly weighted. In Bayesian inference, for example, the weight at each data point is proportional to its likelihood. We illustrate this idea in the following remark.
Remark 2.1.
If there is a change of measure due to, for example, observed data with likelihood $\pi(y|x)$, then the resulting new empirical measure is given by
$\hat{\mu}_{\rm em}({\rm d}x)=C^{-1}\sum_{i=1}^{M}\pi(y|x^{(i)})\delta_{x^{(i)}}({\rm d}x),$
where $C=\sum_{i=1}^{M}\pi(y|x^{(i)})$. Let us introduce the probability vector
$\hat{p}^{\ast}=C^{-1}(\pi(y|x^{(1)}),\ldots,\pi(y|x^{(M)}))^{\rm T}\in\mathbb{R}^{M}.$
The associated coupling is now provided by
$\hat{P}_{\epsilon}=D(\hat{v}_{\epsilon})T_{\epsilon}D(\hat{v}_{\epsilon})$
subject to
$\hat{P}_{\epsilon}{\rm 1}_{M}=\hat{p}^{\ast}.$
Furthermore, the transition probabilities $p_{j}\in\mathbb{R}^{M}$, $j=1,\ldots,M$, get replaced by
$\hat{p}_{j}=\frac{\hat{P}_{\epsilon}e_{j}}{1_{M}^{\rm T}\hat{P}_{\epsilon}e_{j}}\in\mathbb{R}^{M}$
and the associated Markov chain will sample from the empirical posterior distribution $\hat{\mu}_{\rm em}({\rm d}x)$.
@@
3. Diffusion approximation
In order to implement (3), we need to define $m_{\epsilon}(x)$ and $\Sigma(x)$ for $x\in\mathbb{R}^{d}$. In this section, we discuss how one can obtain these functions from the training samples $\{x^{(i)}\}_{i=1}^{M}$ and the diffusion map approximation (6).
Given the underlying reference process (4) and samples $\{x^{(i)}\}_{i=1}^{M}$, recall that $X^{\prime}|X=x\sim N(x,\epsilon(K(x)+K(X^{\prime})))$. We therefore introduce the vector $t_{\epsilon}(x)\in\mathbb{R}^{M}$ with entries
(7) $t_{\epsilon,i}(x)=\exp\left(-\frac{1}{2\epsilon}(x^{(i)}-x)^{\rm T}\left(K(x)+K(x^{(i)})\right)^{-1}(x^{(i)}-x)\right)$

### v2

# Stable generative modeling using Schrödinger bridges

@@ Abstract.
We consider the problem of sampling from an unknown distribution for which only a sufficiently large number of training samples are available. Such settings have recently drawn considerable interest in the context of generative modelling and Bayesian inference. In this paper, we propose a generative model combining Schrödinger bridges and Langevin dynamics. Schrödinger bridges over an appropriate reversible reference process are used to approximate the conditional transition probability from the available training samples, which is then implemented in a discrete-time reversible Langevin sampler to generate new samples. By setting the kernel bandwidth in the reference process to match the time step size used in the unadjusted Langevin algorithm, our method effectively circumvents any stability issues typically associated with the time-stepping of stiff stochastic differential equations. Moreover, we introduce a novel split-step scheme, ensuring that the generated samples remain within the convex hull of the training samples. Our framework can be naturally extended to generate conditional samples and to Bayesian inference problems. We demonstrate the performance of our proposed scheme through experiments on synthetic datasets with increasing dimensions and on a stochastic subgrid-scale parametrization conditional sampling problem.
Keywords: Langevin dynamics, Schrödinger bridges, generative modeling, Bayesian inference
AMS: 60H10,62F15,62F30,65C05,65C40
@@ 1. Introduction
Generative modeling is the process of learning a mechanism for synthesizing new samples that resemble those of the original data-generating distribution, given only a finite set of samples. It has seen wide adoption and enormous success across diverse application domains, from image [5, 22] and text generation [55, 31], to drug discovery [3, 2] and anomaly detection [10, 47], to name but a few.
In this paper, we introduce a new nonparametric approach to generative modeling that combines ideas from Schrödinger bridges [41, 6] and reversible Langevin dynamics [40]. Suppose that we are given $M$ training samples $x^{(i)}\sim\pi$, $i=1,\ldots,M$, from an unknown distribution $\pi$ on $\mathbb{R}^{d}$. Perhaps the simplest nonparametric approach to generative modeling is to build a kernel density estimate (KDE) and then sample from it; the KDE is essentially a mixture model with $M$ components. Alternatively, one could estimate the score function, $s(x)\approx\nabla\log\pi(x)$, without directly estimating $\pi$, and use this estimate as the drift term of Langevin dynamics,
(1) $\dot{X}_{\tau}=s(X_{\tau})+\sqrt{2}\dot{W}_{\tau},$
where $W_{\tau}$ denotes standard $d$-dimensional Brownian motion. There is a plethora of ways of estimating the score function [24, 51], and given an estimate for it, one needs to discretize (1), for example using Euler–Maruyama, to obtain an implementable scheme. However, the step size needs to be carefully chosen: a small step size leads to slow convergence, while too large a step yields instability of the numerical scheme, especially for data that are supported on a compact manifold, e.g.,
@@
for some unknown function $g(x)$. Any estimated score function $\hat{s}_{M}(x)$ will take large values for $x$ with $\|g(x)\|^{2}\gg 0$, rendering the Langevin dynamics (1) stiff. This implies that an explicit time integration method such as Euler–Maruyama will require very small step sizes $\Delta\tau>0$.
In this paper, we choose an alternative approach. Instead of first estimating the score function $\hat{s}_{M}(x)$ and then discretising (1) in time, we employ Schrödinger bridges [41, 6] to directly estimate the conditional expectation value
(3) $\mu_{\epsilon}(x):=\mathbb{E}[X_{\epsilon}|X_{0}=x]$
from the given samples $\{x^{(i)}\}_{i=1}^{M}$ for given parameter $\epsilon>0$. We denote this approximation by $m_{\epsilon}(x):\mathbb{R}^{d}\to\mathbb{R}^{d}$. The second key ingredient of our method is to interpret $\epsilon$ as a step size and to read off a Gaussian transition kernel from the Schrödinger bridge approximation in the form of
(4) $X_{n+1}=m_{\epsilon}(X_{n})+\sqrt{\Sigma(X_{n})}\Xi_{n},$
with appropriately defined diffusion matrix $\Sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ and $\Xi_{n}\sim{\rm N}(0,\epsilon I)$. Broadly, $m_{\epsilon}(X_{n})$ controls the drift, while $\sqrt{\Sigma(X_{n})}\Xi_{n}$ introduces noise. An obvious choice for $\Sigma(x)$ is $\Sigma(x)=2I$, which corresponds to (1) and its Euler–Maruyama discretisation with step-size $\Delta\tau=\epsilon$.
@@
In the recursion (4), the step size $\epsilon$ is linked to the error of the Schrödinger bridges approximation. Comparing to directly discretizing (1) using Euler–Maruyama with step-size $\Delta\tau=\epsilon$, we will demonstrate that the scheme (4) is stable and ergodic for all step-sizes $\epsilon>0$ and, hence, $\epsilon$ can be chosen solely on accuracy considerations. Furthermore, the inclusion of the position-dependent diffusion matrix $\Sigma(x)$ makes it better suited for sampling from a manifold.
@@
Our construction utilizes a Sinkhorn scheme [34, 54], which solves the discrete Schrödinger bridge problem of optimally coupling the empirical measure of the training samples with itself over an appropriate diffusion process. By solving the Schrödinger bridge problem, we construct a transition matrix whose state space encompasses all the training points. We generalize this approach to the continuous state space that extends beyond the current training data points. Armed with this transition kernel, which directly leads to the desired approximation $m_{\epsilon}(x)$, we obtain a Markov chain that samples from the underlying distribution of the training data via (4).
From here, we introduce a novel split-step time-stepping scheme, which ensures that the generated samples consistently lie within the convex hull of the training samples and is ergodic [40]. In contrast, using a direct discretization of (1) with Euler–Maruyama results in generating samples on unbounded domains.
In addition, we replace the constant diffusion matrix $\Sigma(x)=2I$ with a scaled matrix
(5) $\Sigma(x)=2\rho(x)I$
@@
for given bandwidth $\rho(x)>0$, which requires appropriate modifications to the Schrödinger bridges considered in [54]. As we demonstrate in our numerical experiments, the resulting sampling scheme (4) provides a better representation of the underlying target distribution. More precisely, we assess the quality of the generated samples using a variable bandwidth kernel and using a fixed bandwidth kernel on synthetic data sets drawn from non-uniform distributions supported on irregular domains and on low-dimensional manifolds. In addition, we explore data-informed choices of $\Sigma(x)$ in (4).
@@
We then extend our method to cover Bayesian inference problems with $\pi$ as prior and to create a conditional generative model. These extensions allow us to perform Bayesian inference in the “simulation-based” setting, i.e., without explicit evaluation of a prior density and, in the case of conditional sampling, even without evaluations of the data likelihood. We demonstrate the performance of our conditional generative model for a stochastic subgrid-scale parametrization problem.
We close this introduction by pointing to a close connection of the considered Schrödinger bridge problems to diffusion maps [8, 7, 39] which have also been employed to approximate the semigroup $\exp(\epsilon\mathcal{L})$, $\epsilon>0$, of a diffusion process with generator $\mathcal{L}$ and invariant distribution $\pi$, from given samples $\{x^{(i)}\}_{i=1}^{M}$. We recall that $\mu_{\epsilon}(x)=\exp(\epsilon\mathcal{L}){\rm Id}(x)$, where ${\rm Id}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the identity map. However, as demonstrated in [54], the Schrödinger bridge approach leads to a more accurate approximation $m_{\epsilon}(x)$ of the analytic $\mu_{\epsilon}(x)$ in terms of $\epsilon>0$.
@@ 1.1. Related work
Langevin dynamics (1) characterizes the motion of particles as they experience a blend of deterministic and stochastic forces. Unlike in this paper, it is typically assumed that the deterministic forcing term $\nabla\log\pi(x)$ is given. Langevin dynamics has been used as a popular tool for sampling data from the target distribution $\pi$. One variation of this is to introduce a symmetric preconditioning operator to the Langevin dynamics and to consider reversible processes of the general form
(6) $\dot{X}_{\tau}=K(X_{\tau})\nabla\log\pi(X_{\tau})+\nabla\cdot K(X_{\tau})+\sqrt{2K(X_{\tau})}\dot{W}_{\tau},$
where $K(x)$ is a symmetric positive definite matrix and we rely on the Itô interpretation of the multiplicative noise term [40]. While (6) has originally been discussed in molecular physics [12, 11, 23], popular choices of $K(x)$ arising from computational statistics include the empirical covariance [13] and the Riemannian metric [14, 30], making this method converge faster and more geometry-aware, while leaving the stationary distribution unchanged. The scaled diffusion matrix (5) corresponds to the choice $K(x)=\rho(x)I$. Optimal choice of $K(x)$ in terms of convergence to equilibrium have recently been discussed in [29].
@@
Our approximation of the semigroup $\exp(\epsilon\mathcal{L})$ relies on recent work on diffusion maps [34, 54], which relies on an accelerated Sinkhorn algorithm. The Sinkhorn algorithm solves for the Markov transition kernel associated with a discrete Schrödinger bridge problem, where the coupling is between the empirical measure of the training samples with itself. This approach results in a symmetric stochastic operator that, notably, approximates the semigroup $\exp(\epsilon\mathcal{L})$ to higher accuracy in $\epsilon$ than standard diffusion map approximations [8, 7]. Separately, the idea of using variable bandwidth kernels can be found very early in the statistics community, for example, in the context of kernel density estimation [46, 52]. Recently, [1] replaces the original fixed bandwidth kernel with the variable bandwidth kernel in the construction of diffusion maps, making the approximation of the generator accurate on unbounded domains. Inspired by this concept, we replace the fixed bandwidth kernel $\Sigma=2I$ with a variable bandwidth kernel (5) for given $\rho(x)>0$. The resulting Schrödinger bridge approximates the semigroup of the reversible Langevin diffusion process (6) with $K(x)=\rho(x)I$.
@@
In recent years, there has been a surge of research interest in the realm of generative modeling. Despite the remarkable achievements of well-established neural network-based generative models, such as variational auto-encoders (VAE) [26, 44], generative adversarial networks (GAN) [15], and diffusion models or score generative models (SGM) [21, 49], they often require meticulous hyperparameter tuning [45, 50] and exhibit a long training time [15, 53]. Furthermore, SGMs solve both a forward and a reverse stochastic differential equation (SDE). The forward SDE introduces noise to the sample, transforming the data into the standard normal distribution, while the reverse SDE takes sample from the standard normal distribution back to the original data distribution, yielding a different sample than the one initially fed into the forward SDE. During the training process, the score function is learned, not for the target distribution, but for the data distribution at each time. Our work, on the other hand, solves only one (forward) SDE, and we learn the score of the target distribution only once.
@@
Several recent studies have combined a range of score function estimation techniques with Langevin dynamics. For example, [49] introduces a noise conditional score network to learn the score function and then uses annealed Langevin dynamics to generate samples. [4] studies the convergence rate of a Langevin based generative model, where the score is estimated using denoising auto-encoders. Such techniques are also studied within the Bayesian imaging community, commonly referred to as “plug and play” [28]. These approaches use neural networks for the estimation of the score function, necessitating substantial fine-tuning, and their effectiveness depends on factors such as the complexity of the approximation families and the architectural structures of the neural networks. In addition, [28] uses an explicit projection to ensure that samples stay on the compact manifold given by (2) which is assumed to be explicitly known. In contrast, we do not assume any knowledge of $\mathcal{M}$.
1.2. Outline
@@
In Section 2 we construct a Markov chain using a Schrödinger bridge approximation that samples from the given discrete data distribution. In Section 3, we extend this Markov chain to the continuous state space setting by constructing a Gaussian transition kernel which extracts its conditional mean and covariance matrix from the underlying diffusion map approximation. We introduce two discrete-time Langevin samplers; one with an arbitrary data-unaware diffusion and one with a data-aware diffusion matrix in Section 3.1. Theoretical properties such as stability and ergodicity are discussed in Section 3.2. We further discuss the application of variable bandwidth kernels when constructing the Schrödinger bridge in Section 3.3. While Section 3 focuses on finite step-size and finite sample-size implementations, Section 4 establishes connections to the underlying semi-groups and generators in the infinite sample-size limit. We explore the extension of our proposed scheme to a conditional sampling setting and Bayesian inference in Section 5. We demonstrate our proposed methods in Section 6 in a suite of examples, including a conditional sampling exercise with an application to stochastic subgrid-scale parametrization. We conclude in Section 7 with a summary and an outlook.
@@ 2. Discrete Schrödinger bridges
We consider the Schrödinger bridge problem of coupling $\mu_{\rm em}$ with itself along a reversible reference process with (unnormalized) transition probabilities
(7) $t_{ij}=\exp\left(-\frac{1}{2\epsilon}(x^{(i)}-x^{(j)})^{\top}\left(K(x^{(i)})+K(x^{(j)})\right)^{-1}(x^{(i)}-x^{(j)})\right),$
which we collect into a symmetric matrix $T_{\epsilon}\in\mathbb{R}^{M\times M}$. Here $\epsilon>0$ is a tuneable parameter and $K(x)$ is a symmetric positive definite matrix for all $x\in\mathbb{R}^{d}$. Popular choices include $K=I$, $K=\Sigma_{M}$, where $\Sigma_{M}$ is the empirical covariance matrix of the samples $\{x^{(i)}\}_{i=1}^{M}$, and $K=\rho(x)I$, where $\rho(x)>0$ is a scaling function representing variable bandwidth.
Instead of working with the empirical measure $\mu_{\rm em}({\rm d}x)$, we introduce the probability vector $p^{\ast}=(1/M,\ldots,1/M)^{\top}\in\mathbb{R}^{M}$ over $\{x^{(i)}\}_{i=1}^{M}$. Then the associated Schrödinger bridge problem can be reformulated into finding the non-negative scaling vector $v_{\epsilon}\in\mathbb{R}^{M}$ such that the symmetric matrix
@@
is a Markov chain with invariant distribution $p^{\ast}$, i.e.,
$P_{\epsilon}p^{\ast}=p^{\ast}.$
Here $D(v)\in\mathbb{R}^{M\times M}$ denotes the diagonal matrix with diagonal entries provided by $v\in\mathbb{R}^{M}$. We remark that the standard scaling used in Schrödinger bridges would lead to a bistochastic matrix $\tilde{P}_{\epsilon}$, which is related to (8) by $\tilde{P}_{\epsilon}=M^{-1}P_{\epsilon}$.
Given $P_{\epsilon}$, one can now construct a Monte Carlo scheme that samples from $\mu_{\rm em}$. Assume the Markov chain is currently in state $x^{(j)}$, then the transition probabilities to the next state $x\in\{x^{(i)}\}_{i=1}^{M}$ are given by
$p_{j}=P_{\epsilon}e_{j}\in\mathbb{R}^{M},$
where $e_{j}\in\mathbb{R}^{M}$ denotes the $j$-th unit vector in $\mathbb{R}^{M}$. Since all entries in $P_{\epsilon}$ are bounded from below provided all samples satisfy $x^{(i)}\in\mathcal{M}$, where $\mathcal{M}$ is a compact submanifold in $\mathbb{R}^{d}$, the constructed Monte Carlo scheme possesses a unique invariant measure given by $p^{\ast}$ and is geometrically ergodic. The rate of convergence can be determined by the diffusion distance
@@
If the diffusion distance is small, then $x^{(i)}$ and $x^{(j)}$ are well connected. Furthermore, if $d(x^{(i)},x^{(j)})$ is small for all points, then the Markov chain will mix quickly. In particular, larger values of $\epsilon$ will lead to faster mixing.
@@
3. Approximating the conditional mean
In order to implement (4), we need to define $m_{\epsilon}(x)$ and $\Sigma(x)$ for $x\in\mathbb{R}^{d}$. In this section, we discuss how one can obtain these functions from the training samples $\{x^{(i)}\}_{i=1}^{M}$ and the Markov chain approximation (8).
We introduce the vector $t_{\epsilon}(x)\in\mathbb{R}^{M}$ with entries
(9) $t_{\epsilon,i}(x)=\exp\left(-\frac{1}{2\epsilon}(x^{(i)}-x)^{\top}\left(K(x)+K(x^{(i)})\right)^{-1}(x^{(i)}-x)\right)$